There were 340 people in a picnic. There were 20 more men than women and 20
more children than adults.
i) Write the above problem in the form of system of linear equations
iii) How many men were there in the picnic? Use Cramer’s Rule to calculate
your answer.
matriculation mathematics

Answer by rapture1965(18)

Let m = men
Let w = women
Let c = children
m = 20 + x
w = x
c = 20 + (20 + x) + x

20 + x + x + 20 + 20 + 2x = 340
4x + 60 = 340
4x = 340 - 60
4x = 280
x = 280/4
x = 70
There are 70 women.
We know that men = 20 + x.
There are 20 + 70 or 90 men.
We know there are more children than adults.
c = 20 + (20 + 70) + 70
c = 20 + 90 + 70
c = 90 + 90
c = 180
There are 180 children.
Understand?